OpenStudy (anonymous):

slope of 2/3 and passes through 1,0

OpenStudy (candyme):

Your formula is y=mx+b Start with that and fill in what you know. y=2/3x+b Now plug in tthe point and solve for b. 0=2/3(1)+b 0=2/3+b b=-2/3 Plug that back in y=2/3x-2/3